# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ V Discussion
Table XV illustrates the performance of different scenarios, highlighting variations in network traffic and the significance of a unified approach using the M-En dataset. In our experiments, we first trained the model with IoT data tested it with traditional (trad.) data, and then trained it on trad. data and tested it with IoT data. The models performed poorly in both cases due to the distinct traffic patterns, demonstrating that existing SOTA datasets are inadequate for creating a unified framework for M-En networks. Subsequently, we conducted experiments using the M-En dataset, training models on it, and testing with both IoT and trad. traffic. The models showed significant improvements in handling both types of traffic, underscoring the effectiveness of our M-En dataset for training models in diverse scenarios.
TABLE XV: Results using Different Training and Testing Datasets
Training Testing Accuracy Class Precision Recall F1-Score
IoT Trad. 0.0066 0 0.01 0.03 0.01
@@
Trad. IoT 0.0142 0 0.13 0.02 0.03
1 0.00 0.00 0.00
M-En IoT and Trad. 0.9993 0 1.00 1.00 1.00
1 1.00 1.00 1.00
Furthermore, Figure 10 demonstrates the performance of our testing in a real-time scenario. Figure 10(a), illustrates the connection between the server and the gateway of M-En networks, which facilitates the real-time ingestion of network packets for processing. Once the connection is established, the models begin analyzing the traffic. Initially, the performance is suboptimal, showing accuracy levels around 50-60%. However, as the system continues to operate, the performance progressively improves due to the continuous learning approach implemented in our framework. The continuous learning mechanism enables the models to adapt to the evolving traffic patterns by retraining on newly collected data in real-time as shown in Figure 10(b). Over time, the accuracy significantly increases, stabilizing at 99-100%, as seen in the later stages of the testing process. This improvement highlights the robustness and adaptability of our framework, as well as the importance of incorporating real-time data and continual learning for addressing the dynamic nature of M-En network environments. These results further validate the effectiveness of our unified framework and its ability to handle diverse and evolving traffic scenarios in real-time applications.
@@
Significance: This study has several significance in network security for DDoS attack detection and shows strong contribution in the given domain:
  • This study introduces a comprehensive benchmark dataset specifically tailored for M-En networks, providing a foundational resource for researchers to advance security solutions in complex, M-En network settings.
  • The availability of our research resources in an open repository not only ensures transparency and reproducibility but also encourages further development, validation, and deployment of novel methodologies in the domain of M-En network security.
  • By incorporating continuous learning and human intervention, the framework demonstrates an ability to adapt dynamically to new and unseen traffic patterns. This adaptability is essential for dealing with zero-day attacks and evolving threats, providing a sustainable approach to long-term network security.
  • Multi-LF ability to leverage a lightweight model (M1) for initial detection significantly reduces computational overhead, which is crucial for real-time applications. This ensures that the framework is not only accurate but also efficient, making it suitable for deployment in resource-constrained environments.
Limitations: Despite the significance of this study, there are several limitations to consider:
  • This study focuses primarily on a few DDoS attacks and does not explore other types of attacks, such as Man-in-the-Middle (MiM) and other sophisticated network attacks, which restricts the comprehensiveness of the dataset.
  • This study focuses solely on IoT and traditional networks within the M-En framework, which is limited. Numerous other networks, including SDN and industrial IoT, can be incorporated into M-En networks, which could be helpful in bigger networks.
  • Multi-LF requries human intervention, which can slow down the decision-making process. On average, a human expert takes approximately 1 to 2 minutes to make a decision. To address this, a faster and more efficient processing system is needed for timely decision-making.
VI Conclusion
This study effectively addresses DDoS attack detection in M-En networks using a continuous learning approach. A pivotal contribution of this research is the creation of a novel benchmark dataset specifically designed for M-En networks, integrating both IoT and traditional IP-based traffic patterns. Utilizing this comprehensive dataset, we developed a highly reliable methodology optimized for real-time operational scenarios. Our extensive analysis demonstrates that the M-En network dataset exhibits distinct characteristics compared to existing domain-specific benchmark datasets, such as greater variability in packet size and frequency. In contrast, domain-specific datasets tend to have consistent traffic patterns. Consequently, ML models trained on domain-specific datasets underperform when applied to diverse traffic domains, highlighting the inherent challenges in developing a unified detection framework. However, the inherent diversity within the M-En dataset enables the development of a unified system that can reduce computational costs, resource usage, and management overhead. Furthermore, the study concludes that models trained without a continuous learning paradigm suffer significant performance degradation over time due to evolving traffic patterns and dynamic network topologies, such as the addition or removal of devices. In contrast, our proposed Multi-LF, incorporating continuous learning mechanisms, effectively mitigates these issues by enabling the model to adapt to zero-day attacks and scale efficiently within M-En networks. This results in enhanced reliability and sustained accuracy over time. Additionally, Multi-LF achieves high classification accuracy and low prediction latency in real-time testing environments, although offline testing exhibits marginally superior performance metrics. Importantly, Multi-LF continuously improves its real-time performance by fine-tuning with newly acquired data, thereby maintaining robustness against previously unseen attack vectors. This study also concludes that human involvement during the continuous learning process is highly effective in avoiding bias and improving transparency and trust in the framework.
However, despite the significant contributions of this study, we acknowledge certain limitations and outline directions for future work. We intend to explore additional unique attack types to further enhance the system’s efficiency. Additionally, we consider integrating Large Language Models (LLMs) to replace human intervention, enabling faster and more reliable decision-making with increased trust.
@@ Acknowledgment
This work is funded by the School of Computer Science and CHIST-ERA ERA-NET - SPiDDS Topic and Irish Research Council (IRC) for funding support
Author Contributions
Furqan Rustam: Writing – original draft, Visualization, Software, Validation, Methodology, Investigation, Formal analysis, Data curation, Conceptualization, Writing – review & editing. Islam Obaidat: Visualization, Validation, Investigation, Formal analysis, Data Curation, Writing – review & editing. Anca Delia Jurcut: Writing – review & editing, Validation, Supervision, Software, Project administration, Formal analysis, Conceptualization.
Appendix A Biography Section
Furqan Rustam is currently pursuing a Ph.D. degree in Computer Science from University College Dublin, Ireland. He received his MCS degree in the Department of Computer Science, Islamia University of Bahawalpur, Pakistan (Oct-2015 to Oct-2017). In Nov-2018, he got himself enrolled in the Master of Computer Science, Department of Computer Science, Khwaja Fareed University of Engineering and Information Technology (KFUEIT), Rahim Yar Khan, 64200, Pakistan. He also served as a Research Assistant at Fareed Computing & Research Center, KFUEIT, Pakistan. His recent research interests are related to Data Mining, Machine Learning, and Artificial Intelligence, mainly working on Creative Computing and Supervised Machine Learning.
Islam Obaidat an Assistant Professor at North Carolina A&T State University, specializing in cybersecurity, IoT software security, and computer networks. He earned his Ph.D. in Computer Science and his research includes developing testbeds for simulating IoT botnet DDoS attacks, enhancing DDoS attack simulations through full system emulation, and creating large-scale memory error IoT botnets using NS3DockerEmulator. Dr. Obaidat’s work aims to advance the understanding and mitigation of security threats in IoT environments.
ANCA D. JURCUT received a Bachelor of Mathematics and Computer Science degree from the West University of Timisoara, Romania, in 2007, and a Ph.D. degree from the University of Limerick, Ireland, in 2013. From 2008 to 2013, she was a Research Assistant with the Data Communication Security Laboratory, University of Limerick. From 2013 to 2015, she was working as a Postdoctoral Researcher with the Department of Electronic and Computer Engineering, University of Limerick. She was a Software Engineer with IBM, Ireland. Since 2015, she has been an Assistant Professor with the School of Computer Science, University College Dublin (UCD), Ireland. She is currently an Assistant Professor at the School of Computer Science, UCD. Her research interests include network and data security, security for the Internet of Things (IoT), security protocols, formal verification techniques, and applications of blockchain technologies in cybersecurity.

### v2

@@ 5 Discussion
To justify the need for a robust AI solution for M-En networks and the limitations of existing datasets, we conducted cross-domain experiments. Table 20 shows performance variations across scenarios, underscoring the value of a unified M-En-based approach. In our experiments, we first trained the model with IoT data, tested it with traditional (trad.) data, and then trained it on trad. data and tested it with IoT data. The models performed poorly in both cases due to the distinct traffic patterns, demonstrating that existing SOTA datasets are inadequate for creating a unified framework for M-En networks. Subsequently, we conducted experiments using the M-En dataset, training models on it, and testing with both IoT and trad. traffic. The models showed significant improvements in handling both types of traffic, underscoring the effectiveness of our M-En dataset for training models in diverse scenarios.
Table 20: Results Using Different Training & Testing Datasets
Training Testing Acc Class Pre Rec F1
[…]
Trad. IoT 0.014 0 0.13 0.02 0.03
@@
[…]
Furthermore, Figure 10 demonstrates the performance of our testing in a real-time scenario. Figure 10(a) illustrates the connection between the server and the gateway of M-En networks, which facilitates the real-time ingestion of network packets for processing. Once the connection is established, the models begin analyzing the traffic. Initially, the performance is suboptimal, showing accuracy levels around 50-60%. However, as the system continues to operate, the performance progressively improves due to the continuous learning approach implemented in our framework. The continuous learning mechanism enables the models to adapt to the evolving traffic patterns by retraining on newly collected data in real-time, as shown in Figure 10(b). Over time, the accuracy significantly increases, stabilizing at 99-100%, as seen in the later stages of the testing process. This improvement highlights the robustness and adaptability of our framework, as well as the importance of incorporating real-time data and continual learning for addressing the dynamic nature of M-En network environments. These results further validate the effectiveness of our unified framework and its ability to handle diverse and evolving traffic scenarios in real-time applications.
@@
$\Rightarrow$ Significance: This study has several significances in network security for malicious attack detection and shows a strong contribution in the given domain: (i) This study introduces a comprehensive benchmark dataset specifically tailored for M-En networks, providing a foundational resource for researchers to advance security solutions in complex, M-En network settings. (ii) The availability of our research resources in an open repository not only ensures transparency and reproducibility but also encourages further development, validation, and deployment of novel methodologies in the domain of M-En network security. (iii) By incorporating continuous learning and human intervention, the framework demonstrates an ability to adapt dynamically to new and unseen traffic patterns. This adaptability is essential for dealing with zero-day attacks and evolving threats, providing a sustainable approach to long-term network security. (iv) Multi-LFability to leverage a lightweight model (M1) for initial detection significantly reduces computational overhead, which is crucial for real-time applications. This workflow ensures that the framework is not only accurate but also efficient, making it suitable for deployment in resource-constrained environments.
$\Rightarrow$ Limitations: Despite the significance of this study, there are several limitations to consider: (i) The present work focuses on IoT and traditional networks within the M-En framework. This controlled scope ensures methodological clarity and reproducibility. However, the framework is inherently extensible and can seamlessly integrate additional domains, such as SDN and consumer electronics (CE) networks, during real-time deployment, enabling broader applicability in larger and more heterogeneous infrastructures. (ii) Multi-LFincorporates human oversight to validate critical or uncertain detections. While this may introduce latency under high-alert conditions, it enhances interpretability and operational reliability, which are key for security in sensitive environments. Notably, studies and industry reports indicate that a human analyst typically manages 100–200 alerts per day [^5] https://countershadow.com/blog/ai-vs-cyber-threats-accelerating-response-in-a-24-7-landscape/, underscoring the value of incorporating guided automation to reduce analyst workload. Future work will focus on integrating active learning and feedback-driven tuning to streamline this process further. (iii) The continual learning strategy retrains the model using previously misclassified samples, enhancing adaptability to evolving threats. Although this iterative update process increases computational demand over time, it ensures the model remains robust against concept drift. Future optimization, such as lightweight retraining or selective memory replay, can maintain adaptability while minimizing resource overhead.
6 Conclusion & Future Direction
In this paper, we first constructed a packet-level M-En dataset (from live DDoShield-IoT traffic mixed with public PCAP traces) to provide the research community with a realistic corpus and a reproducible pipeline for heterogeneous M-En attack research. Building on this foundation, we presented Multi-LF, a continuous-learning framework that unifies lightweight and deep architecture models and incorporates selective human feedback to secure M-En networks. We deployed Multi-LF online in DDoShield-IoT  using a Transformer-based M1 & M2, maintaining 0.999 accuracy while requiring human intervention to label only 0.0026% for 115,795 live packets. In contrast, the baseline approach requires significantly more human intervention, 109 expert checks out of 111,644 packets, which represents 0.0976% of the workload. This online evaluation confirms a high detection rate under real traffic, while the chosen transformer models provide high throughput and consume low power, resulting in 0.007 kg CO2 emission per experiment. We concluded that traditional models suffer a drop in accuracy during online testing, whereas our Multi-LFeffectively handles live traffic through layered verification and minimal human intervention, achieving near error-free performance. Additionally, our statistical feature extraction configuration, using a 1-second processing interval, proved effective. Features like average payload size and destination port entropy significantly improved malicious traffic detection. Further, our findings show that continuous learning with weight interpolation enables the M1 model to adapt to new traffic and attacks while preserving prior knowledge, minimizing human effort.
In future work, we aim to expand the dataset with encrypted and zero-trust traffic and explore unsupervised drift detection to enable adaptive retraining. We also plan to enrich DDoShield-IoT  by injecting additional binaries and diverse attack types, along with augmenting it using public PCAP  repositories via our modular pipeline. To evaluate realism, we will compute entropy and protocol-distribution metrics and benchmark them against backbone traces (e.g., MAWI [cho2000traffic]). Finally, we intend to explore the integration of Large Language Models (LLMs) to support or automate human-in-the-loop decisions for improved speed and reliability.
Data Availability Statement
The code and datasets used for reproducibility of the results in this study are publicly available on GitHub at https://github.com/furqanrustam/MultiLF.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Furqan Rustam: Conceptualization, Data curation, Formal analysis, Validation, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft. Islam Obaidat: Data curation, Formal analysis, Investigation, Validation, Resources, Writing – review & editing. Anca Delia Jurcut: Formal analysis, Investigation, Project administration, Resources, Supervision, Writing – review & editing.
@@ Acknowledgment
This work is funded by the School of Computer Science, CHIST-ERA ERA-NET - SPiDDS Topic, and the Irish Research Council (IRC).
Appendix A Statistical Feature Definitions and Calculation Details
Statistical Feature 1. Packet Count: The total number of packets observed in each time window, represented as $\text{PacketCount}_{t}=\sum_{i=1}^{n}\text{Packet}_{i}$. 2. Destination Port Entropy: Measures the entropy of destination ports to detect scanning activities, defined as $\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i})$. 3. Most Frequent Source Port: Identifies the most common source port in a window, $\text{SourcePort}_{\max}$. 4. Most Frequent Destination Port: Identifies the most common destination port in a window, $\text{DestinationPort}_{\max}$. 5. Short-lived Connections: Counts the number of short-lived connections, $\text{ShortLivedConnections}=\sum_{i=1}^{n}\delta_{i}$, where $\delta_{i}$ indicates a short-lived connection. 6. Repeated Connection Attempts: Measures repeated connection attempts, $\text{RepeatedAttempts}=\sum_{i=1}^{n}\alpha_{i}$, where $\alpha_{i}$ indicates a repeated attempt. 7. Network Scanning Activity: Counts instances of SYN flags without ACK flags, $\text{Scanning}=\sum_{i=1}^{n}(\text{SYN}-\text{ACK})$. 8. Flow Rate: Calculated as packets per second, $\text{FlowRate}=\frac{\text{TotalPackets}}{\text{TimeInterval}}$. 9. Source Entropy: Entropy of source addresses, $\text{SourceEntropy}=-\sum_{i=1}^{n}q_{i}\log(q_{i})$, where $q_{i}$ is the probability distribution of source addresses. 10. Connection Errors (RST flag): Counts instances of RST flags, $\text{RSTCount}=\sum_{i=1}^{n}\text{RST}_{i}$. 11. Most Frequent Packet Size Frequency: Identifies the most common packet size, $\text{PacketSize}_{\max}$. 12. Abnormal Size Frequency: Counts packets exceeding a size threshold, $\text{AbnormalSizeFrequency}=\sum_{i=1}^{n}\text{Packet}_{i}\text{ if size}_{i}>\text{Threshold}$. 13. Sequence Number Variance: Variance in sequence numbers, $\text{Var}(\text{SequenceNumber})$. 14. Average Packet Number: Average packets per interval, $\text{AvgPackets}=\frac{\text{TotalPackets}}{\text{TimeIntervals}}$. 15. SYN Frequency: Frequency of SYN flags, $\text{SYNFrequency}=\frac{\text{TotalSYN}}{\text{TimeInterval}}$. 16. ACK Frequency: Frequency of ACK flags, $\text{ACKFrequency}=\frac{\text{TotalACK}}{\text{TimeInterval}}$. 17. TCP Frequency: Proportion of TCP packets, $\text{TCPFrequency}=\frac{\text{TotalTCP}}{\text{TotalPackets}}$. 18. UDP Frequency: Proportion of UDP packets, $\text{UDPFrequency}=\frac{\text{TotalUDP}}{\text{TotalPackets}}$. 19. Most Frequent Protocol: Most used protocol, $\text{Protocol}_{\max}$. 20. Packet Size Variability: Variance in packet sizes, $\text{Var}(\text{PacketSize})$. 21. Most Frequent Payload Size: Most common payload size, $\text{PayloadSize}_{\max}$. 22. Average Payload Size: Mean payload size, $\text{AvgPayloadSize}=\frac{\text{TotalPayload}}{\text{TotalPackets}}$. 23. Packet Size Standard Deviation: Standard deviation of packet sizes, $\text{StdDev}(\text{PacketSize})$. 24. Average Packet Size: Mean size of packets within a window, $\text{AvgPacketSize}=\frac{\text{TotalPacketSize}}{\text{TotalPackets}}$.